OpenStudy (anonymous):

six times the reciprocal of a number is equal to 4 times the reciprocal of 2. find the number. How would I write this out?

OpenStudy (anonymous):

6*(1/x) = 4*(1/2); 6/x = 2; 2x = 6; x = 3

OpenStudy (anonymous):

6*(1/x)=4*(1/2)

OpenStudy (anonymous):

6/x=2

OpenStudy (anonymous):

i mean x=6/2

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

be careful, because x was in the denominator, not the numerator

OpenStudy (anonymous):

u beat me again D;

OpenStudy (anonymous):

ackkk sorry XD I will stop as I have physics homework to do anyhow

OpenStudy (anonymous):

ME TOO

OpenStudy (anonymous):

but mine's really easy

OpenStudy (anonymous):

XDDD oh isn't life a joy ;D

OpenStudy (anonymous):

i'll just have to see how mine goes :)

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

np

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

good luck with your physics hw

OpenStudy (anonymous):

would you guys minde helping with one more?

OpenStudy (anonymous):

if it is super quick : D as i have to do my own now lol

OpenStudy (anonymous):

it takes john 4 hrs to mow a lawn it takes maria 3 hrs to mow lawn how long would it take if they worked together?

OpenStudy (anonymous):

t = time required if they work together; (t/3) + (t/4) = 1 completed job; multiply by 12: 4t + 3t = 12; 7t = 12; t = 12/7 = 1.7142857..... hr. I THINK.

OpenStudy (anonymous):

round probably : )

OpenStudy (anonymous):

hope that makes sense

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

np